OpenStudy (anonymous):

Solve for x. -3x+2/5 = 7/1

OpenStudy (ash2326):

we have \[-3x+\frac 25=7\]@Sunshine447 multiply whole equation by 5, what would you get?

Parth (parthkohli):

Similarly, you may also subtract \(2 \over 5\) from both sides. :)

OpenStudy (anonymous):

\[-3x + 2/5 = 7 ................ \times 5\] \[-15x + 2 = 35\] ........... Take 2 to the other side \[-15x = 33\] \[X = -33/15\]

OpenStudy (anonymous):